SW updates and configuration settings.

FL test failed. Roberto started to investigate.

The problem seems to be related to the crate #3 (BCU board we swapped on 20/02).

Swapped #1 with #3 for confirming BCU or backplane problem. cr#2-3 Not working: seems backplane problem.=> cr#0-1 works. Confirmed: problem related to crate #3 itself or connection between #2 and #3. Confirmed daisy chain #2 and #3 FL broken. Removed and susbtituted.

Flux meter calibration done.pbgain matrix 18,2,3 (contando da 1) = 6.5e-3 del PB#3.-> only changed housekeepergain.h file

TSS usage: disabilita tss / accendi i coils / abilita tss

-- MarcoXompero - 21 Feb 2010
Topic revision: r6 - 22 Feb 2010, MarcoXompero
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback